ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 185, 2011

Adam, D. The Gel′fond–Fridman theorem in positive characteristic  235

Balan, R., P. Casazza and Z. Landau Redundancy for localized frames

445

Baro, E. On the o-minimal LS-category  61

Bates, T., S. Eilers and D. Pask Reducibility of covers of AFT shifts

207

Beiglböck, M. An ultrafilter approach to Jin’s Theorem  369
Bergner, J. E. Homotopy fiber products of homotopy theories  389
Brofferio, S. and B. Schapira Poisson boundary of GL_{d}(ℚ)  125

Casazza, P. See R. Balan

Caselli, F. Projective reflection groups  155

Costantini, M. and G. Zacher On the subgroup index problem in finite groups

293

Eilers, S. See T. Bates Fel, L. G. Duality relation for the Hilbert series of almost symmetric

numerical semigroups

413

Johnson, W. B. and B. Zheng Subspaces and quotients of Banach spaces with shrinking unconditional bases

375

Kaplan, I., T. Scanlon and F. O. Wagner Artin–Schreier extensions in NIP and simple fields

141

Koldobsky, A. Positive definite functions and stable random vectors  277

Landau, Z. See R. Balan Meyerovitch, T. Quasi-factors and relative entropy for infinite-

measure-preserving transformations

43

Moravec, P. On the Schur multipliers of finite p-groups of given coclass

189

Panchishkin, A. Families of Siegel modular forms, L-functions and modularity lifting conjectures

343

Pask, D. See T. Bates Patassini, M. On the irreducibility of the Dirichlet polynomial of a

simple group of Lie type

477

Scanlon, T. See I. Kaplan Schapira, B. See S. Brofferio Stepanov, A. and N. Vavilov On the length of commutators in

Chevalley groups

253

Temkin, M. Relative Riemann–Zariski spaces  1

Vavilov, N. See A. Stepanov Wagner, F. O. See I. Kaplan Warshall, A. D. A group with deep pockets for all finite generating

sets

317

Yamana, S. Degenerate principal series representations for quaternionic unitary groups

77

Zacher, G. See M. Costantini Zheng, B. See W. B. Johnson